# VGC-Bench: A Benchmark for Generalizing Across Diverse Team Strategies in Competitive Pokémon

### v2

@@ 4.2 Baseline Implementations
4.2.1 Trained Baselines
@@
All the policy networks used across the learning baselines embed each agent’s moves, items, and abilities into latent representations, and then use an additional aggregation token and a 3-layer Transformer encoder [23] to aggregate the information of 12 Pokémon (6 from the ego teams and 6 from the opponent’s team) in the field. When frame stacking is enabled, we use a second Transformer encoder along the time axis, apply a positional encoding, and use causal masking to process the historical information and only use the logits from the last frame. The output of the Transformer encoder is then projected to the size of the action space by a linear layer into the logits before the softmax layer. The logits at the invalid actions are masked by . We also handle interdependent action constraints, such as ensuring that both Pokémon do not switch into the same replacement. For reinforcement learning methods, we adopt an actor-critic implementation without parameter sharing, though they share the same network architecture.
Behavior Cloning.
We train a reward-filtered behavior cloning (BC) policy [6] to match the distribution of human actions given a state from the dataset collected by the human play data collector. Specifically, we select trajectories from replays in which players have ladder ratings greater than or equal to , ensuring high-quality demonstrations. The parameterized policy is trained to imitate human decision-making by minimizing the negative log-likelihood of the demonstrated actions:
Population-based Reinforcement Learning.
We employ an actor-critic implementation of the Proximal Policy Optimization (PPO) [18] to train all reinforcement learning agents. For agents to approach the desired minimax solution of the game via RL, we implement three population-based training paradigms from the unified empirical game-theoretic framework [11] on top of RL: self-play, fictitious play, and double oracle [13]. In self-play (SP), the agent trains against itself. The fictitious play (FP) variant maintains a pool of the agent’s past checkpoints, and the agent learns against a uniform distribution of its past selves, resetting the opponent policies each episode. We take the final response policy to the policy pool as the method’s output. Double oracle (DO) differs from FP in that it derives the Nash equilibrium distribution from a maintained empirical payoff matrix between all agents in the pool, and uses that distribution to sample opponents. We solve the Nash distribution by solving a linear programming problem because of the game’s two-player, zero-sum, and symmetric-payoff nature. Future work could employ more complex meta-game solvers with the Policy Space Response Oracle [11].
Our preliminary experiments suggest that this population-based framework can potentially perform better when initialized with strong policies. Thus, we also implement baselines that fine-tune the policy from the BC-trained parameters (BCSP, BCFP, BCDO). Each method is trained with the same number of training teams and interaction timesteps. The hyperparameters for RL (details in Table 10) were tuned to serve as a reasonable starting point for future work.
4.2.2 Other Baselines
@@
We also provide a basic LLM agent that uses Meta-Llama-3.1-8B-Instruct [7]. Each turn, the battle state is converted into a prompt that is then fed into the agent along with an available action space, instructions on how to construct a valid action, and the required format of its response. If the LLM agent does not respond in the correct format, the agent selects a random action. Note that the LLM agent was not thoroughly researched; we acknowledge that it is possible to make a stronger LLM agent with tool use and other techniques [10]. However, this LLM agent baseline provides a reasonable starting point on which research can be conducted.
@@
From poke-env, we inherit 3 rule-based baselines: RandomPlayer, MaxBasePowerPlayer, and SimpleHeuristicsPlayer [17]. The RandomPlayer plays random moves every turn, the MaxBasePowerPlayer greedily chooses the highest-power move from the available options every turn, and the SimpleHeuristicsPlayer has a hard-coded weighted sum of multiple heuristics to select the best move based on the heuristic scores, based on information such as current HP of the active Pokémon, type matchup viability, accuracy, and power of its moves, etc. We extend SimpleHeuristicsPlayer so that it can also play VGC battles in a double format.
4.3 Evaluation Methods
VGC-Bench provides multiple methods of evaluation, including cross-evaluation of agents, training agents to exploit a fixed opponent, and a means to compute custom ELO ratings (the details of which are described in Appendix B). The cross-evaluation over games of a pair of policies from the policy pool , where , , and , can be defined point-wise as
where is a set of team configurations that the pair of policies both sample uniformly from each game. The combination of all these evaluation primitives enables insight into agent performance, generalization, and exploitability.
1.
Performance can be tested by cross-playing agents with teams that every agent had experience with during training. Ideally, as you train agents with more teams, performance would minimally decrease on an individual team basis compared with agents trained with fewer teams. We define as the set of teams configurations that policy was trained on; the evaluation set of teams is constrained by
@@ 4.3 Evaluation Methods
2.
Generalization can be tested by cross-playing agents with teams outside of their training distribution; if an agent trained with more teams outperforms agents trained on fewer teams, this suggests that the agent has generalization capabilities. We constrain the evaluation set of teams by
@@
Exploitability can be tested by approximating a best response policy against the agent to be evaluated, , via RL training and measuring the highest win rate that the best response can achieve [22]. This provides a lower bound on the true exploitability of the agent. Whatever maximum win rate the counter-policy achieves over a sufficiently large number of training steps serves as our exploitability measurement for our agent. For an agent trained on teams, a best response policy must be trained for each individual team, and the exploitability of the agent is defined as
(6) In practice, we initialize our counter-policy with the BC agent since a random counter-policy initialization can struggle to get a reward signal against a strong agent. If the best response policy does not achieve a 100% win rate against the agent being tested, the user may run multiple trials to tighten the lower bound of the highest possible win rate that is being measured.
5 Experiments
@@
To encourage research in Pokémon VGC, we gather preliminary results interrogating the following questions: (1) Is it possible to use standard RL methods to create an AI agent that plays at a human expert level for the one-team setting? (2) How do performance, generalization, and exploitability scale as the training team size increases? All experiments are conducted on a cluster with 8 A40 GPUs and 2 Intel(R) Xeon(R) Gold 6342 CPU @ 2.80GHz.
5.1 Cross-Play Performance
@@
For all agents that fine-tuned with RL, we did training runs with team set sizes of 1, 3, 10, and 30. For each team set size, we compare our 6 RL baseline agents, the BC agent, 3 rule-based players, and our LLM agent (all described in Section 4.2.1) through cross-play in Figure 3. Each win rate entry of the cross-play matrix is calculated with episodes with a uniformly drawn team for each game played. For the actual win count values, see Tables 5, 6, 7, 8 in the Appendix C.
Notice that in the one-team setting (whose win counts are recorded in table 5), FP beats BCFP 83 to 17, BCFP beats DO 83 to 17, and DO beats FP 62 to 38 – evidence that Pokémon VGC is a cyclic game. However, the heatmaps become more transitive as the training team size increases.
We also calculate ELO ratings based on the payoff matrices for each team set size as shown in Table 2. Note that these ELO ratings are calculated within the VGC-Bench benchmark, not on the online pokemon-showdown ladder. The online ladder is not useful for evaluating performance on in-distribution teams since the opponent’s team configuration cannot be controlled, but we acknowledge that the ladder is a useful metric to measure out-of-distribution performance. It is clear from Figure 3 and Table 2 that while non-BC methods achieve decent performance, they are generally outperformed by methods initialized with BC. It is not clear from these results which of self-play, fictitious play, and double oracle are most effective, but fictitious play seems to be the most reliably high-performance.
| R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO | |
| 1 | 1127 | 1520 | 1621 | 1204 | 1681 | 1717 | 1573 | 1463 | 1721 | 1768 | 1627 |
| 3 | 1122 | 1443 | 1585 | 1174 | 1631 | 1609 | 1523 | 1370 | 1736 | 1689 | 1708 |
| 10 | 1124 | 1466 | 1602 | 1170 | 1634 | 1604 | 1562 | 1444 | 1692 | 1730 | 1743 |
| 30 | 1052 | 1398 | 1544 | 1069 | 1508 | 1538 | 1514 | 1361 | 1612 | 1647 | 1642 |
5.2 Performance and Generalization Evaluations
@@
We then measure performance and generalization as described in Section 4.3. We used the highest overall ELO-rated agents in the 1, 3, 10, and 30-team settings for this evaluation; Table 4 shows that as the size of the set of teams used during training increases, the performance of the AI agent on any one team decreases considerably and consistently. These results indicate that our agent does not pass the performance test even for relatively small , highlighting the central challenge that Pokémon VGC poses to the AI community. To test generalization, we used 4 out-of-distribution teams and compared agents across team set sizes via cross-play. In Table 4, we can see that our agent exhibits generalization since win rate increases moderately as increases.
| #Teams | 1 | 3 | 10 | 30 |
| 1 | – | 79 | 83 | 92 |
| 3 | 21 | – | 68 | 82 |
| 10 | 17 | 32 | – | 60 |
| 30 | 8 | 18 | 40 | – |
| #Teams | 1 | 3 | 10 | 30 |
| 1 | – | 55 | 34 | 33 |
| 3 | 45 | – | 30 | 33 |
| 10 | 66 | 70 | – | 55 |
| 30 | 67 | 67 | 45 | – |
5.3 Exploitability Evaluation
We measured the exploitability of the strongest ELO-rated agent from each of our 4 team set sizes as described in Section 4.3. For each, we initialized the exploiter agent with the BC agent’s parameters. We found that for all 4 agents, after 3 million training steps, the exploiter agent was able to achieve approximately 100% win rate.
5.4 Human Evaluation
@@
After measuring ELO ratings and evaluating performance across team set sizes, we tested our most performant agent in the most performant team set size – the behavior cloning agent fine-tuned with fictitious play in the 1-team setting – against an intermediate, advanced, and expert player. Against the intermediate player, our agent won all 5 matches, and against the advanced player, the agent won 2 out of 5 total matches played. Indeed, our agent was even able to win matches against the expert-level player who has qualified for the VGC World Championships. To the best of our knowledge, our agent is the first AI system to achieve such a feat. We received feedback that although the agent is very strong on initial play, it does have noticeable dips in performance in certain states. After enough successive games, strong human players can adapt and beat the agent.
6 Conclusion
@@
In this work, we introduced VGC-Bench, a benchmark designed to evaluate the generalization capabilities of AI agents in the challenging and combinatorially complex environment of Pokémon VGC. Our benchmark includes standardized evaluation protocols, a modular training pipeline, curated human gameplay data, and implementations of a broad set of baseline methods ranging from behavior cloning and reinforcement learning to game-theoretic algorithms like fictitious play and double oracle. We also contributed substantial improvements to the poke-env library, including integration with PettingZoo and environment support for VGC formats, thereby enabling easy adoption by the broader research community.
@@
Looking forward, we identify two major research directions enabled by VGC-Bench:
1.
Generalization to Teams, : Our current agents achieve expert-level performance in the single-team setting but struggle as the number of evaluated teams increases. A natural extension is to develop agents that can generalize across team matchups, evaluate metagames through cross-play matrices, and ultimately perform at a superhuman level regardless of the team piloted. Researchers can use the experiments in Tables 4 and 4 to assess progress on this front.
2.

### v3

@@ 4.2. Baseline Implementations
4.2.1. Trained Baselines
Behavior Cloning.
We train a behavior cloning (BC) policy (gleave2022imitation) to match the distribution of human actions given a state from the dataset collected by the human play data collector, without the use of our rating or winner-based data filters. The parameterized policy is trained to imitate human decision-making by minimizing the negative log-likelihood of the demonstrated actions:
Multi-Agent Reinforcement Learning.
We employ an actor-critic implementation of Proximal Policy Optimization (PPO) (schulman2017ppo) to train all reinforcement learning agents. For agents to approach the desired minimax solution of the game via RL, we implement three multi-agent training paradigms from the unified empirical game-theoretic framework (lanctot2017unified) on top of RL: self-play, fictitious play, and double oracle (mcmahan2003planning). In self-play (SP), the agent trains against itself. The fictitious play (FP) variant maintains a pool of the agent’s past checkpoints, and the agent learns against a uniform distribution of its past selves, resampling the opponent policies after every experience-gathering period. We take the final response policy to the policy pool as the method’s output. Double oracle (DO) differs from FP in that it derives the Nash equilibrium distribution from a maintained empirical payoff matrix between all agents in the pool, and uses that distribution to sample opponents. We solve the Nash distribution by solving a linear programming problem because of the game’s two-player, zero-sum, and symmetric-payoff nature. Future work could employ more complex meta-game solvers with the Policy Space Response Oracle (lanctot2017unified). We also implement baselines that fine-tune the policy from the BC-trained parameters (BCSP, BCFP, BCDO). Each method is trained with the same number of training teams and interaction timesteps. The hyperparameters for RL used for all PPO methods (details in Table 7) were tuned to serve as a reasonable starting point for future work.
@@ 4.2.1. Trained Baselines
All the policy networks used across the learning baselines embed each agent’s moves, items, and abilities into latent representations, and then use an additional aggregation token and a 3-layer Transformer encoder (vaswani2017attention) to aggregate the information of 12 Pokémon (6 from the agent’s team and 6 from the opponent’s team) in the field. When frame stacking is enabled, we use a second Transformer encoder along the time axis, apply a positional encoding, and use causal masking to process the historical information and only use the logits from the last frame. The output of the Transformer encoder is then projected to the size of the action space by a linear layer into the logits before the softmax layer. The logits at the invalid actions are masked by . We also handle interdependent action constraints, such as ensuring that both Pokémon do not switch into the same replacement. For reinforcement learning methods, we adopt an actor-critic implementation without parameter sharing, though they share the same network architecture.
4.2.2. Other Baselines
@@
We also provide a basic LLM agent that uses Meta-Llama-3.1-8B-Instruct (grattafiori2024llama). Each turn, the battle state is converted into a prompt that is then fed into the agent along with an available action space, instructions on how to construct a valid action, and the required format of its response. If the LLM agent does not respond in the correct format, the agent selects a random action. Note that the LLM agent was not thoroughly researched; we acknowledge that it is possible to make a stronger LLM agent with tool use and other techniques (karten2025pok). However, this LLM agent baseline provides a reasonable starting point on which research can be conducted.
@@
From poke-env, we inherit 3 rule-based baselines: RandomPlayer, MaxBasePowerPlayer, and SimpleHeuristicsPlayer (poke_env). The RandomPlayer plays random moves every turn, the MaxBasePowerPlayer greedily chooses the highest-power move from the available options every turn, and the SimpleHeuristicsPlayer has a hard-coded weighted sum of multiple heuristics to select the best move based on the heuristic scores, which considers information such as current HP of the active Pokémon, type matchup viability, accuracy, power of its moves, etc. We extend SimpleHeuristicsPlayer so that it can also play double battles for the VGC format.
4.3. Evaluation Methods
VGC-Bench provides multiple methods of evaluation, including cross-evaluation of agents, Alpha-Rank (omidshafiei2019alpha) for ranking agents based on their cross-evaluation performance, training an exploiter policy against an agent to lower-bound the agent’s worst-case lose-rate, and evaluating agents over varying team set sizes with the performance and generalization tests.
Let be the set of team configurations used for evaluation and be the uniform distribution of those teams. The cross-evaluation of a pair of policies from the policy pool , where , , and , can be defined as
where controls team , controls team , and the distribution of trajectories is defined in Equation 1. Alpha-Rank (omidshafiei2019alpha) provides a way of transforming this cross-play matrix into an ordered ranking of policies, allowing us to determine a best policy among a pool.
The combination of these evaluation primitives enables insight into agent performance, generalization, and exploitability.
Performance can be tested by cross-playing agents with teams that every agent had experience with during training. Ideally, as you train agents with more teams, performance would minimally decrease on an individual team basis against agents trained with fewer teams. We define as the set of teams configurations that policy was trained on; the evaluation set of teams is constrained by
@@ 4.3. Evaluation Methods
Generalization can be tested by cross-playing agents with teams that none of them experienced during training; how much an agent trained on more team configurations surpasses those trained on fewer configurations reflects its generalization ability. We constrain the evaluation set of teams by
@@
Exploitability can be tested by approximating a best response policy, , against the agent to be evaluated via RL training and measuring the highest win rate that the best response can achieve (timbers2020approximate). Whatever maximum win rate the exploiter agent achieves over a sufficiently large number of training steps serves as our exploitability measurement for our agent. We formally define the exploitability of the agent as
5. Experiments
@@
To encourage research in Pokémon VGC, we gather preliminary results interrogating 2 main research questions. In sections 5.2 and 5.3, we investigate how performance, generalization, and exploitability scale as the training team size increases. In section 5.4 we see if it is possible to use standard RL methods to create an AI agent that plays at a human expert level for the one-team setting. All experiments are conducted on a cluster with 8 A40 GPUs and 2 Intel(R) Xeon(R) Gold 6342 CPU @ 2.80GHz.
5.1. Cross-Play Performance
Cross-Play Win Rate Heatmaps for varying team set sizes.
@@ 5.1. Cross-Play Performance
For all agents that fine-tuned with RL, we did training runs with team set sizes of 1, 4, 16, and 64. For each team set size, we compare our 6 RL baseline agents, the BC agent, 3 rule-based players, and our LLM agent (all described in section 4.2) through cross-play in Figure 3. Each win-rate entry of the cross-play matrix is calculated with 1000 episodes with uniformly drawn teams for each game played, except for the LLM player, which was evaluated with 100 episodes due to the LLM’s relatively slow inference. These 1000 episodes equally evaluate across 5 training seeds for all learning methods. For each seed, a different set of teams is selected for the 1, 4, 16, and 64-team settings, and these team sets are always perfectly nested so that, for example, all teams from the 16-team setting of a given seed are included in the 64-team setting of that seed. All cross-evaluations have win-rates relative to the row players, with errors of approximately for 1000-game evaluations, and for the LLM player’s 100-game evaluations. For numerical win rates, see Tables 8, 9, 10, 11 in Appendix B.
Notice that as the number of teams increases, the heatmaps become progressively less noisy. We hypothesize that this is likely caused by (1) all of our learning methods noticeably degrading in performance for larger team set sizes, and (2) results being averaged over larger numbers of matchups reducing matchup-specific biases in the evaluation.
| 1 Team | 4 Teams | 16 Teams | 64 Teams | |
| R | 11 | 9 | 11 | 11 |
| MBP | 9 | 7 | 9 | 9 |
| SH | 7 | 5 | 8 | 7 |
| LLM | 10 | 8 | 10 | 10 |
| SP | 4 | 4 | 6 | 4 |
| FP | 6 | 4 | 5 | 8 |
| DO | 5 | 4 | 4 | 5 |
| BC | 8 | 6 | 7 | 6 |
| BCSP | 1 | 1 | 3 | 1 |
| BCFP | 2 | 2 | 2 | 3 |
| BCDO | 3 | 3 | 1 | 2 |
This table contains a relative ranking of each baseline agent featured in VGC-Bench. Each agent is only compared with other agents that trained with the same number of teams.
It is clear from Figure 3 and Table 2 that while non-BC methods achieve decent performance, they are generally outperformed by methods initialized with BC. Self play seems to be the most reliably high-performance method in the 1 team setting, but we note that the performance differences between the BC methods for the 4, 16, and 64 team set sizes are negligible.
5.2. Performance and Generalization Evaluations
@@
We then measure performance and generalization as described in section 4.3. We used the highest overall ranked agents from our Alpha-Rank evaluation in the 1, 4, 16, and 64-team settings for this evaluation; Table 3 shows that as the size of the set of teams used during training increases, the performance of the AI agent on any one team decreases considerably and consistently. These results indicate that our agent does not pass the performance test even for relatively small , highlighting the central challenge that Pokémon VGC poses to the AI community. To test generalization, we used 72 out-of-distribution teams and compared agents across team set sizes via cross-play. In Table 4, we can see that our agent exhibits generalization since win rate increases moderately as increases.
| #Teams | 1 (BCSP) | 4 (BCSP) | 16 (BCDO) | 64 (BCSP) |
| 1 (BCSP) | – | 0.699 | 0.74 | 0.698 |
| 4 (BCSP | 0.301 | – | 0.594 | 0.672 |
| 16 (BCDO) | 0.26 | 0.406 | – | 0.644 |
| 64 (BCSP) | 0.302 | 0.328 | 0.356 | – |
| #Teams | 1 (BCSP) | 4 (BCSP) | 16 (BCDO) | 64 (BCSP) |
| 1 (BCSP | – | 0.405 | 0.375 | 0.331 |
| 4 (BCSP) | 0.595 | – | 0.453 | 0.422 |
| 16 (BCDO) | 0.625 | 0.547 | – | 0.436 |
| 64 (BCSP) | 0.669 | 0.578 | 0.564 | – |
In order to solidify the validity of our generalization results, we also calculate statistics regarding the similarity of the 72 teams we set aside during training with the upper bound of 64 teams used during training in Table 5. We measure team similarity on a scale of 0 to 1, taking into account matching species between teams, and among those matching species, whether they have the same item, moves, ability, Tera type, and stat configurations. We ensure that no two teams are fully identical, and assert that the average similarity being about 0.5 across all training seeds firmly grounds the Generalization Test as a legitimate experiment.
| Run | Mean | Median | Min | Max |
| 1 | 0.508 | 0.513 | 0.268 | 0.946 |
| 2 | 0.522 | 0.526 | 0.221 | 0.947 |
| 3 | 0.530 | 0.524 | 0.221 | 0.948 |
| 4 | 0.557 | 0.554 | 0.260 | 0.947 |
| 5 | 0.526 | 0.526 | 0.238 | 0.948 |
5.3. Exploitability Evaluation
We measured the exploitability of the strongest agent according to the Alpha-Rank evaluation from each of our team set sizes as described in section 4.3. This evaluation was averaged across the agents from all 5 training seeds, and all exploiter agents were trained on the same team configuration from the 1-team setting, regardless of how many teams the agent being exploited experienced during training. We tested exploitability by initializing the exploiter with a random initialization in Table 4 and with the pre-trained BC policy in Table 5. In almost all cases, all agents are approximately 100% exploitable, although the best agent in the 1-team setting does exhibit notably stronger resistance to exploitation when the exploiter agent is randomly initialized.
An exploiter agent initialized randomly trains to exploit the strongest agent as determined in Table 2.
An exploiter agent initialized as the behavior cloning agent trains to exploit the strongest agent as determined in Table 2.
5.4. Human Evaluation
@@
After measuring Alpha-Rank ratings and evaluating performance across team set sizes, we tested our most performant agent in the most performant team set size – the behavior cloning agent fine-tuned with self-play in the 1-team setting – against an intermediate, advanced, and expert player. Our expert player was Aaron Traylor, a multi-time competitor in the World Championships and the 2020 Dallas Regional Champion. All 3 human testers were instructed to try to find a way to exploit the AI system over 5 consecutive games and encouraged to think for as long as they needed to. Against the intermediate player, our agent won all 5 matches, and against the advanced player, the agent won 2 out of 5 total matches played. Indeed, our agent was even able to win against the expert-level player. To the best of our knowledge, our agent is the first AI system to achieve such a feat. We received feedback that although the agent is strong on initial play, it does have noticeable dips in performance in certain states. After enough successive games, strong human players can adapt and beat the agent. Please note that these human evaluations are intentionally anecdotal; more comprehensive human evaluation should be conducted before reaching stronger conclusions based on these results.
6. Conclusion and Future Work
@@
In this work, we introduced VGC-Bench, a benchmark designed to evaluate the generalization capabilities of AI agents in the challenging and combinatorially complex environment of Pokémon VGC. Our benchmark includes standardized evaluation protocols, a modular training pipeline, curated human gameplay data, and implementations of a broad set of baseline methods ranging from behavior cloning and reinforcement learning with game-theoretic algorithms like self-play, fictitious play, and double oracle, the strongest of which was able to win against a past World Championships competitor. We also contributed substantial improvements to the Poké-env library, including integration with PettingZoo and environment support for VGC formats, thereby enabling easy adoption by the broader research community.
@@
Looking forward, we identify five major research directions enabled by VGC-Bench:
Generalization to Teams, : Our current agents achieve strong performance in the single-team setting but struggle as the number of teams in training increases. A natural extension is to develop agents that can generalize across team matchups without having performance degrade, and ultimately perform at a superhuman level across arbitrary teams without needing to retrain. Researchers can use the experiments in Tables 3 and 4 to assess progress on this front.
@@
Search: With an accurate and fast simulator, one could model future positions of the game to make more well-informed decisions (brown2020combining; sokota2025superhuman). The benefits of model-based techniques could be bountiful in the domain of Pokémon VGC, as it gives the agent a way to understand how the individual actions of each active Pokémon on the field contribute to the resulting state. In fact, we suspect that much of the difficulty of achieving strong performance across a large set of teams with model-free RL may be that the complexity of predicting future outcomes of the game simply can’t fit into the deep neural network. Past work in Pokémon suggests that search has significant potential (huang2019self; wang2024winning; karten2025pok), but active work is ongoing to create a highly accurate, fast, and open-source simulator.
Opponent Modeling: One could incorporate explicit opponent modeling to infer latent properties of the opposing strategy, such as playstyle, risk tolerance, and long-term tactical intent. By predicting how an opponent is likely to act – or which strategic archetype they belong to – an agent can dynamically adapt its policy rather than treating all opponents as identically optimal or uniformly random (he2016opponent; yu2022model). It may be a good idea to start the opponent model out as a behavior cloning agent from human data and attempt to converge to the opponent’s style as the game plays out, effectively adapting to the opponent in real time from a reasonable starting guess.
Latent Variable Behavior Cloning: Future work could use latent variable imitation learning, enabling the behavior cloning agent to model diverse human play-styles rather than collapsing behavior into a single averaged policy, which could improve its strength by making its play more cohesive and consistent (wang2017robust; hsiao2019learning).
This work has taken place in the Learning Agents Research Group (LARG) at UT Austin. LARG research is supported in part by NSF (FAIN-2019844, NRT-2125858), ONR (W911NF-25-1-0065), ARO (W911NF-23-2-0004), DARPA (Cooperative Agreement HR00112520004 on Ad Hoc Teamwork) Lockheed Martin, and UT Austin’s Good Systems grand challenge. Peter Stone serves as the Chief Scientist of Sony AI and receives financial compensation for that role. The terms of this arrangement have been reviewed and approved by the University of Texas at Austin in accordance with its policy on objectivity in research.
We are very thankful to our human testers: Dylan Remillard (intermediate), Jiaheng Hu (advanced), and Aaron Traylor (expert). Aaron Traylor was also kind enough to explain some niche mechanics in Pokémon VGC, give helpful feedback to the AI’s strengths and weaknesses, connect us with other professional VGC competitors, and even used their platform to help popularize our work. Also, we thank our anonymous reviewers for their diligence in evaluating our work and for the constructive criticism which ultimately helped this paper become what it is.